# Arxiv preprint

### v1

@@ 6 Preliminary Findings
We apply StoryScore to the stories generated by our use-case pipeline (Section 3) on the full test set, in order to assess whether the metric behaves consistently beyond a small illustrative subset.
Table 3 reports aggregated statistics across the test set (76 generated stories) comparing two versions of the pipeline, using either a pre-trained version of Qwen2.5, or a fine-tuned version of the model. Fine-tuning substantially improves StoryScore and Context Recall, and completely eliminates prompt leakage, while other components show more moderate gains. Prompt Cleanliness and Title Coverage saturated on the test set, but their presence in StoryScore is a diagnostic safeguards. No Redundancy and No Hallucination exhibit higher variance, reflecting the sensitivity of long-form generation to repetition artifacts and entity-level unsupported mentions. Nevertheless, the pre-trained model still attains positive scores, particularly on BERTScore, Title Coverage, No Redundancy, and No Hallucination.
@@
We conducted a manual inspection and consistently found that narratives generated in the pre-trained settings are affected by prompt leakage, redundancy, and excessive narrative filler, deficiencies that are only weakly penalized by global semantic metrics. BERTScore operates at the embedding level and is tolerant of paraphrasing, repetition, and generic formulations, allowing long and semantically “safe” texts to achieve high scores despite being poorly readable or editorially unacceptable. Conversely, the fine-tuned model generates more controlled, dense, and readable narratives, yet these qualitative improvements are only partially reflected in metrics. Among the components, Context Recall aligns most clearly with qualitative analysis. Its relatively low values reflect the task’s emphasis on synthesis and reformulation rather than lexical reuse; lower scores for the pre-trained model are consistent with its greater reliance on generic fillers that weaken lexical grounding. Overall, variations in StoryScore were consistent with qualitative judgments of readability, narrative control, and factual grounding.
@@ 7 Conclusions and Future Work
This paper addressed the problem of evaluating AI-generated scientific stories, a task that combines factual grounding with narrative transformation and persona adaptation. We introduced StoryScore, a composite metric that integrates semantic alignment, lexical grounding, structural fidelity, fluency, and hallucination control, and analysed its behaviour on a concrete scientific storytelling use case. Our analysis shows that different metrics capture complementary but uneven aspects of narrative quality. While embedding-based measures reliably assess semantic similarity, they remain largely insensitive to discourse-level control issues such as redundancy, prompt leakage, and narrative structure. The combination of complementary metrics provides a more informative picture of system behaviour than any single score alone. Overall, the initial evaluation suggests that the automatic metrics composing StoryScore are useful for comparisons between generated stories, even if insufficient for discriminating critical aspects of narrative quality and text control in an absolute way.
Hallucination detection proved to be an intrinsically complex task, even more when creativity and pedagogical reformulation are involved. Existing hallucination detection approaches are either too shallow, too rigid, or too unstable. Future evaluation frameworks should explicitly account for persona adaptation and narrative transformation, calling for a new definition of hallucination that distinguish between acceptable abstraction and factual distortion. Finally, this work motivates further refinement of composite metrics such as StoryScore – particularly through the inclusion of components explicitly targeting narrative control and structural validity – as well as an assessment of alignment with human judgement following some best practices [chhun-etal-2022-human].

### v2

@@ 5 Preliminary Findings
We apply StoryScore to a set of 76 stories generated by our use case pipeline. Table 3 reports aggregated statistics comparing two versions of the pipeline, using either a pre-trained version of Qwen2.5, or a fine-tuned version of the model. Fine-tuning substantially improves StoryScore and completely eliminates prompt leakage. Prompt Cleanliness and Title Coverage saturated, but their presence in StoryScore is a diagnostic safeguards. Nevertheless, the pre-trained model still attains positive scores, particularly on BERTScore, Title Coverage, No Redundancy, and No Hallucination.
@@
A manual inspection found that narratives generated in the pre-trained settings are consistently affected by prompt leakage, redundancy, and excessive narrative filler, deficiencies that are only weakly penalized by global semantic metrics. The latters are tolerant of paraphrasing, repetition, and generic formulations, rewarding long and semantically “safe” texts despite being poorly readable. Conversely, the fine-tuned model generates more controlled, dense, and readable narratives, yet these qualitative improvements are only partially reflected in metrics. Overall, variations in StoryScore were consistent with qualitative judgments of readability, narrative control, and factual grounding.
@@ 6 Conclusions and Future Work
We introduced StoryScore, a composite metric that integrates semantic alignment, lexical grounding, structural fidelity, fluency, and hallucination control, and analysed its behaviour on a concrete scientific storytelling use case. The combination of complementary metrics provides a more informative picture of system behaviour than any single score alone. The initial evaluation suggests that StoryScore is useful for comparisons between generated stories, even if insufficient for discriminating critical aspects of narrative quality and text control in an absolute way.
Hallucination detection proved to be an intrinsically complex task, even more when creativity and pedagogical reformulation are involved. Existing hallucination detection approaches are either too shallow, too rigid, or too unstable. Future evaluation frameworks should explicitly account for persona adaptation and narrative transformation, calling for a new definition of hallucination that distinguish between acceptable abstraction and factual distortion. Finally, this work motivates further refinement of composite metrics such as StoryScore – in particular to address narrative control and structural validity – as well as an assessment of alignment with human judgement following some best practices [chhun-etal-2022-human].
Acknowledgements.
This work was supported by the French Public Investment Bank (Bpifrance) i-Demo program within the LettRAGraph project (Grant ID DOS0256163/00).